OpenStudy (anonymous):

What is the area of a circle with a radius of 2 meters? Use twenty-two sevenths for pi

OpenStudy (mathstudent55):

\( A = \pi r^2\) \(A = \dfrac{22}{7} (2~m)^2 \) Now simplify.